Chapter 15, Problem 056 GO A Carnot engine has an efficiency of 0.38. The Kelvin temperature of its hot reservoir is quadruplhelp please
Answer #1

Let the tempete aturce of hot reservoit is Tik and cold reservoir is Tak efficiency is = 1- 7 2 2 -> 0138 = - 12 1-0.38 = 0.6